Please help me with this problem:
How many gallons of a 70% antifreeze solution must be mixed with 70 gallons of 25% antifreeze to get a mixture that is 60% antifreeze? Use the six step method

Answer by stanbon(75887)

How many gallons of a 70% antifreeze solution must be mixed with 70 gallons of 25% antifreeze to get a mixture that is 60% antifreeze? Use the six step method
---------------------------------------------
Equation:
alcohol + alcohol = alcohol
0.70x + 0.25*70 = 0.60(x+70)
---------------------------------
Multiply thru by 100 to get:
70x + 25*70 = 60x + 60*70
10x = 35*70
x = 35*7
x = 245 gallons (amt of 70% solution needed in the mixture)
=======================
Cheers,
Stan H.